OpenStudy (anonymous):

Assume that a 21-year old men's heart rates at rest are normally distributed with a mean of 71 beats per minute and a standard deviation of 6 beats per minute. If 200 men are examined, how many would you expect to have a heart rate of less then 74?

OpenStudy (anonymous):

is it (200-71)/(6)?

OpenStudy (anonymous):

it comes out to be 21.5 and it says to type a whole number.

OpenStudy (anonymous):

anyone? :/

OpenStudy (anonymous):

x ̅=(∑x)/n x ̅=(∑(x∙f) )/n s=√((∑(x-x ̅ )^2 )/((n-1) )) z=(x-μ)/σ

OpenStudy (anonymous):

idk what one lmao!